OpenStudy (anonymous):

What is the area of a circle that has a diameter of 56 feet? Use 22/7 for p.

OpenStudy (pfenn1):

A=(pi) r^2 where r=radius=D/2 where D=diameter of 56 feet radius =D/2=56/2=28 A=(22/7) (28)^2